# Data-Driven Dynamic Factor Modeling via Manifold Learning

### v1

@@ 5.2 Historical Backtesting
To validate our model with data, we conduct several historical backtests covering major periods of financial crisis and set our scenarios equal to the actual realized historical scenarios, such that for a given factor $i\in\mathcal{S}$ in our scenario, the magnitude of the stress would be $\Delta\mathbf{X}_{t}^{i}=\mathbf{X}_{t}^{i}-\mathbf{X}_{t-1}^{i}$. For each of the historical backtests, we implement JDKF and the comparison benchmarks of SSA, Static PCA, and Dynamic PCA, described in Section 5.1. We then compare the conditional (on the scenario) predicted portfolio returns of JDKF, SSA, Static PCA, and Dynamic PCA relative to the realized time $t$ portfolio return. On each period of our backtest, we consider a PM who has access to a universe of $N$ stocks and constructs his portfolio weights via the $1/N$ equal weighting rule. On each period then, the PM predicts the portfolio return conditional on the scenario under each model.
@@
Historical backtesting involves estimating the performance of a given model if it had been employed in the past. Historical backtesting is very well suited for evaluating the performance of scenario analysis models, especially during historical periods of financial stress. The main idea is to attempt to predict portfolio return conditional on a scenario that we have observed in the past. More concretely, if at a given time period $t$ we know the historical magnitude of the stress $\Delta\mathbf{X}_{t}^{i}$ for our scenario-defining factors $i\in\mathcal{S}$, can we accurately predict the co-movements in the un-stressed factors $i\not\in\mathcal{S}$? Of course, the portfolio return is a function of the factors and so from these we compute the conditional scenario portfolio return. Hence the goal of our historical backtest is to assess how well JDKF captures the dynamic correlation among the factors that are driving a portfolio’s overall return as compared to the other benchmarks. Our historical backtesting algorithm, which we describe in detail below, calls on the already defined algorithms in Section 5.1 for each model and treats them as function calls.
Figure 5: Flow of the historical rolling backtest: periodic re-fitting & rolling training window
Our historical backtest, Algorithm 6 below is composed of two phases, the refitting period for periodic re-fitting of the Kalman filters in JDKF and Dynamic PCA, and the backtesting period for the rolling one-step-ahead predictions of portfolio return. Figure 5 displays a schematic representation of our backtesting design. First, for JDKF and Dynamic PCA, we periodically refit the Kalman filters based on a fixed refitting period $R$, for the rolling window training, we select a fixed window size of $s$. The idea is to start the Kalman filters with lead time $R$ and run them through the end of the backtesting period. During the backtesting period, we retrieve the necessary inputs already estimated from the Kalman filters, primarily the filtered estimates for the diffusion coordinates and PC components, and train our models on a one-step-ahead rolling basis, training on data from $t-s:t$ and predicting portfolio return one test period out.
The reason our backtesting design has an additional phase for refitting the Kalman filters, is so that we can avoid refitting the Kalman filter every time the training window rolls one period ahead. The rationale is the well-known burn-in period phenomenon observed in the Kalman filter. The burn-in period refers to the initial phase of the Kalman filter algorithm during which the estimates might not yet accurately reflect the system state. Allowing the filter to run through the burn-in period allows the algorithm to converge to a more stable set of state estimates. For this very reason, we start up the filter with a lead before the backtesting period and let it run for some time in order for the estimates to stabilize.
1:$s,T,K,R$
$\triangleright$ $s=\#$ periods in rolling window for model training
$\triangleright$ $T=\#$ periods in the data
$\triangleright$ $K=\#$ Monte-Carlo samples used to estimate factor model-based scenario Portfolio Return
$\triangleright$ $R=\#$ months before re-fitting the Kalman filter
$\triangleright$ $r$ is the start index for refitting
$\triangleright$ $r_{\text{end}}$ is the end index for refitting
2:for $r\leftarrow 0:R:T-1$ do
3:     $r_{\text{end}}=\min(r+R+s,T)$
6:     for $t\leftarrow\max(s,r+s)$ to $r_{\text{end}}$ do
7:         Retrieve Precomputed Kalman Filter Estimates
8:         Retrieve $\hat{\psi}_{t}$, $\hat{z}_{t}$ from Kalman filter estimates
9:         Estimate Multivariate Regression Model
10:         Fit $\mathbf{\tilde{B}}$ using $(y_{t-s:t},x_{t-s:t})$
11:         Compute Portfolio Simulations
14:         $\hat{V}_{t+1,\text{Static PCA}}\leftarrow\text{StaticPCA\_Simulation}(x_{t-s:t},x_{t+1},\mathbf{\tilde{B}})$……(Algorithm 3)
15:         $\hat{V}_{t+1,\mathrm{SSA}}\leftarrow\text{SSA\_Simulation}(x_{t},x_{t+1},\mathbf{\tilde{B}})$……(Algorithm 2)
16:         Compute true portfolio return $V_{t+1,\mathrm{TRUE}}$
17:         Compute Errors
22:     end for
23:end for
@@
In all of our experiments in Section 6.2, we use $K=10,000$ Monte Carlo simulations of conditional portfolio returns for the probabilistic models JDKF and Dynamic PCA. Since we are working under the special linear multi-factor model case of our JDKF framework, for each refitting period we estimate $\mathbf{B}$ via ordinary least squares (OLS) and $\mathbf{R}$ as the empirical covariance of the OLS residuals; furthermore, we estimate $\mathbf{V}$ as the empirical covariance matrix of the common factors. Given the time series of portfolio return predictions $\hat{V}_{t}$ and absolute errors $E_{t}$ for each model from running the historical backtests, we then consider two main metrics for assessing model performance. Following the literature, as in [haugh2020scenario], we consider the Mean Absolute Error (MAE) for each model defined as:
@@
We also employ standard Value-at-Risk (VaR) exceptions tests to statistically test whether the predicted portfolio returns $\hat{V}_{t,\mathrm{JDKF}}$ from the backtest are consistent with the realized $V_{t,\mathrm{TRUE}}$. As is standard practice in risk management, and using $V_{t}$ as the random variable representing portfolio return, the time $t$ level-$\alpha$ VaR, which is an $\mathcal{F}_{t}$-measurable random variable, is defined for $\alpha\in(0,1)$ as:
@@
$\mathds{1}_{t}(\alpha)=\begin{cases}1&,\text{ if }V_{t}<\operatorname{VaR}_{t}(\alpha)\\ 0&,\text{ otherwise }\end{cases}$
Our backtest gives us a sequence of empirical VaR exception indicators $\hat{\mathds{1}}_{t}(\alpha)$ from the empirical distribution of the JDKF portfolio return predictions. Following [kupiec1995], under the null hypothesis that JDKF is correct, $\sum_{t=1}^{T}\hat{\mathds{1}}_{t}(\alpha)$ follows a Binomial distribution with $T$ trials and success probability $1-\alpha$, from which a two-sided binomial test follows with test statistic:
@@ 6.1 Data Description
We model the series of stock returns via a classical macroeconomic factor model with the 124 FRED-MD macro-level factors from [mccracken2016fred] plus the 8 macroeconomic factors from [goyal2008comprehensive]. This gives us a total of D=132 common factors. We utilize the curated data version as provided from [chen1986economic], which applies some standard transformations to the macro factors to ensure we deal with stationary time series (see Appendix E for the transformation details). The monthly stock returns data is obtained from CRSP. The time period considered is 1967-2016. We consider stocks trading in the NYSE, NASDAQ, and the American Stock Exchange. For each financial crisis period in our backtests, we maximize the size of the universe of firms by using the largest number of firms present throughout the entire time span of the considered period. We select four periods for out-of-sample prediction:
@@
    July 1990 - Mar 1991. We fit the Kalman filters only once with sufficient lead time before the start of the crisis. We re-train for each rolling window time step on all data available from $t-s:t$ and predict conditional portfolio return for each of the 8 periods (months) in the backtesting phase. This period captures to the oil price shock of 1990, debt accumulation, and consumer pessimism present during those times. For this period, we have a universe of N=1,200 firms.
@@
    Mar 2001 - Nov 2001. We fit the Kalman filters only once with sufficient lead time before the start of the crisis. We re-train for each rolling window time step on all data available from $t-s:t$ and predict conditional portfolio return for each of the 8 periods in the backtesting phase. This period captures the dot-com bubble and the 9/11 attacks. For this period we have a universe of N=1,344 firms.
@@
    Dec 2007 - June 2009. We fit the Kalman filters only once, with sufficient lead time before the start of the crisis period. We re-train for each rolling window time step on all data available from $t-s:t$ and predict conditional portfolio return for each of the 18 periods in the backtesting phase. This period contains the subprime mortgage crisis. For this period we have a universe of N=1,409 firms.
@@
For our empirical tests, we consider a portfolio manager who on each period uses the $1/N$ rule to construct the weights of his portfolio. To define our scenarios, we consider the following macroeconomic variables, which roughly correspond to the FED’s 2022 variables for their supervisory scenarios. These variables are taken as our subset $\mathcal{S}$ of the common factors that will define the stress variables in the historical backtests:
Table 3: Scenario Stress Macroeconomic Variables
@@ 6.2 Stress Test Analysis Data
We conduct a historical backtest as outlined in Algorithm 6 for each one of the financial crises. For each backtest, in addition to reporting the performance metrics MAE and percentage accuracy, we also display the time series plots of the predicted conditional portfolio returns for each one of the models along with the true historical portfolio returns, where we re-emphasize that each portfolio is constructed with the $1/N$ rule. An important note is that for Static PCA, we perform each backtest for every possible PC component dimension and select the best performing one in terms of MAE. We begin with the richest of the four backtests, spanning June 2004-December 2016, where we have 150 prediction periods. As mentioned earlier, in this backtest we periodically refit the Kalman filters every 50 months. Figure 6 displays the time series of the conditional predictions from the backtest. We note that for a few months, Dynamic PCA yielded predictions that were significantly far off from the truth, we removed some of these outlier predictions from the plot to preserve the visibility of all the other time series displayed. We display a gray bar over each month where Dynamic PCA yielded such an outlier prediction.
Figure 6: True and conditionally predicted portfolio returns spanning 150 prediction periods. June 2004-December 2016
@@
Table 4: 95% Value at Risk (VaR) Exceptions Results
we also report the results for a 99% VaR exceptions two-sided binomial test:
Metric Value
VaR Threshold ($\alpha=0.01$) -0.1468
Actual Exceptions 1
Expected Exceptions 1.5
Binomial Test p-value 0.99
Table 5: 99% Value at Risk (VaR) Exceptions Results
Note that for both the 95% VaR exceptions test in Table 4 and the 99% VaR exceptions test in Table 5 there is no strong evidence to reject the null hypothesis that JDKF is correct. Meaning that the VaR estimates from JDKF appear statistically consistent with the expected distribution. In both cases the number of observed exceptions is in line with the expected exceptions. For the 95% test we observe 7 exceptions with an expected 7.5, JDKF is slightly conservative but certainly well within an acceptable range. This indicates JDKF is not displaying systematic under or over-estimation of risk.
@@
Figure 7: Diffusion Map Spectrum and Reconstruction Error of the Macroeconomic Factors
@@
Table 6: MAE for JDKF and Benchmark Models Across Different Time Periods
@@
Table 7: Percentage of times JDKF predictions have a higher accuracy, measured by the absolute value error $E_{t}$. We consider four different financial crisis periods.
@@ 7 Conclusions
We combine anisotropic diffusion maps with Kalman filtering to infer the latent dynamic covariate embeddings, and predict the response variable directly from the diffusion map embedding space. We develop a conditional sampling procedure that accounts for dynamic correlations between covariates when conditioning on fixed values for a subset of them, and note that the procedure is general enough that it allows for conditioning on subsets of responses as well. Our approach is justified via theoretical results on the convergence of anisotropic diffusion maps as well as the robustness of approximating the learned diffusion coordinates by linear SDEs. The former result applies concentration inequalities for Markov processes to generalize the arguments of [singer_graph_2006, singer2008nonlinear] to the time series case; the latter adapts arguments typically seen in literature on Markov chain Monte Carlo methods.
@@
While the primary application of our framework in this paper has been financial stress testing, it is important to emphasize the rich set of applications where our framework can be employed. First, our dynamic factor modeling framework is well-suited for asset pricing problems, where one seeks to construct a sparse set of dynamic factors from a high-dimensional factor zoo that possess high explanatory power for the cross-section of stock returns. Another application of our framework is to the problem of nowcasting, where one seeks to utilize a higher-frequency dataset to improve predictions on lower-frequency data. Our data-driven framework gives rise to a broad set of new possibilities, especially when one wishes to make minimal modeling and parametric assumptions on observed data.
[…]

### v2

@@ 6.2 Historical Backtesting
To validate our model with data, we conduct historical backtests covering major periods of financial crisis and set our scenarios equal to the actual realized historical scenarios. For a given factor $i\in\mathcal{S}$ in our scenario, the magnitude of the stress is $\Delta\mathbf{X}_{t}^{i}=\mathbf{X}_{t}^{i}-\mathbf{X}_{t-1}^{i}$. For each historical backtest, we implement JDKF and the comparison benchmarks (SSA, Static PCA, and Dynamic PCA) described in Section 6.1. We then compare the conditional (on the scenario) predicted portfolio returns of each model relative to the realized time $t$ portfolio return. On each period of our backtest, we consider a portfolio manager who constructs portfolio weights via the $1/N$ equal weighting rule and predicts the portfolio return conditional on the scenario under each model.
@@
Historical backtesting involves estimating the performance of a given model if it had been employed in the past. The main idea is to predict portfolio return conditional on a scenario that we have observed in the past. More concretely, if at time period $t$ we know the historical magnitude of the stress $\Delta\mathbf{X}_{t}^{i}$ for our scenario-defining factors $i\in\mathcal{S}$, can we accurately predict the co-movements in the un-stressed factors $i\not\in\mathcal{S}$? The goal of our historical backtest is to assess how well JDKF captures the dynamic correlation among the factors driving a portfolio’s overall return as compared to the other benchmarks. Our backtest consists of two phases (see Figure 5):
  • Refitting Period. For the linear multi-factor JDKF and Dynamic PCA, we periodically refit the Kalman filters based on a fixed refitting period $R$. We start the Kalman filters with lead time $R$ before the backtesting period and run them through to the end, allowing the filters to pass through their burn-in period and reach stable estimates.
  • Backtesting Period. We retrieve the filtered estimates for the diffusion coordinates (for the linear multi-factor JDKF) and PC components (for Dynamic PCA) from the pre-fitted Kalman filters. We then train our models on a one-step-ahead rolling basis using data from ts:$t$-s:t (where $s$ is the rolling window size) and predict portfolio return one period out.
Our rationale for this two-phase design is to avoid refitting the Kalman filter every time the training window rolls forward, which would require the filter to repeatedly pass through its burn-in period. By allowing the filter to run continuously with periodic refitting every $R$ periods, we ensure the estimates remain stable and accurate. Figure 5 displays a schematic representation of our backtesting design.
Figure 5: Structure of the historical rolling backtest with periodic Kalman filter refitting. The timeline is divided into refitting periods and backtesting periods. During each refitting period, the Kalman filter is fitted once to obtain filtered estimates $\{\hat{\psi}_{t}\}$ and $\{\hat{z}_{t}\}$. Within each backtesting period, models are trained on rolling windows and predictions are made one period ahead. We show the full 200-month timeline spanning June 2004-December 2016 with three complete refitting cycles. The insets illustrate that after a burn-in period, the Kalman filter runs through the refitting period to stabilize estimates before rolling predictions begin.
@@
Given the time series of portfolio return predictions $\hat{V}_{t}$ and absolute errors $E_{t}$ for each model from running the historical backtests, we then consider two main metrics for assessing model performance. Following the literature, as in [22], we consider the Mean Absolute Error (MAE) for each model defined as:
@@
We also employ a standard Value-at-Risk (VaR) exceptions test to statistically test whether the predicted portfolio returns $\hat{V}_{t,\mathrm{JDKF}}$ from the backtest are consistent with the realized $V_{t,\mathrm{TRUE}}$. As is standard practice in risk management, and using $V_{t}$ as the random variable representing portfolio return, the time $t$ level-$\alpha$ VaR, which is an $\mathcal{F}_{t}$-measurable random variable, is defined for $\alpha\in(0,1)$ as:
@@
[…]
Our backtest gives us a sequence of empirical VaR exception indicators $\hat{\mathds{1}}_{t}(\alpha)$ from the empirical distribution of the JDKF portfolio return predictions. Following [28], under the null hypothesis that JDKF is correct, $\sum_{t=1}^{T}\hat{\mathds{1}}_{t}(\alpha)$ follows a Binomial distribution with $T$ trials and success probability $1-\alpha$. This yields a two-sided binomial test with test statistic:
@@
The complete implementation details with periodic Kalman filter refitting for the historical rolling backtest are provided in Appendix F.5, including the detailed steps of the procedure stated in Algorithm 6.
@@ 7.1 Data Description
We model the series of stock returns via a classical macroeconomic factor model with the 124 FRED-MD macro-level factors from [30] plus the 8 macroeconomic factors from [21]. This gives us a total of D=132 common factors. We utilize the curated data version as provided from [11], which applies some standard transformations to the macro factors to ensure we deal with stationary time series (see Appendix G for the transformation details). The monthly stock returns data is obtained from CRSP. The time period considered is 1967-2016. We consider stocks trading in the NYSE, NASDAQ, and the American Stock Exchange. For each financial crisis period in our backtests, we maximize the size of the universe of firms by using the largest number of firms present throughout the entire time span of the considered period. We select four periods for out-of-sample prediction:
@@
    July 1990 - Mar 1991. We fit the Kalman filters only once with sufficient lead time before the start of the crisis. We re-train for each rolling window time step on all data available from ts:$t$-s:t and predict conditional portfolio return for each of the 8 periods (months) in the backtesting phase. This period captures to the oil price shock of 1990, debt accumulation, and consumer pessimism present during those times. For this period, we have a universe of N=1,200 firms.
@@
    Mar 2001 - Nov 2001. We fit the Kalman filters only once with sufficient lead time before the start of the crisis. We re-train for each rolling window time step on all data available from ts:$t$-s:t and predict conditional portfolio return for each of the 8 periods in the backtesting phase. This period captures the dot-com bubble and the 9/11 attacks. For this period we have a universe of N=1,344 firms.
@@
    Dec 2007 - June 2009. We fit the Kalman filters only once, with sufficient lead time before the start of the crisis period. We re-train for each rolling window time step on all data available from ts:$t$-s:t and predict conditional portfolio return for each of the 18 periods in the backtesting phase. This period contains the subprime mortgage crisis. For this period we have a universe of N=1,409 firms.
@@
For our empirical tests, we consider a portfolio manager who, in each period, uses the $1/N$ rule to construct the weights of his portfolio. To define our scenarios, we consider the set of macroeconomic variables in Table LABEL:tab:scenario_variables, which roughly correspond to the FED’s 2022 variables for their supervisory scenarios. These variables are taken as our subset $\mathcal{S}$ of the common factors that will define the stress variables in the historical backtests:
Table 3: Scenario Stress Macroeconomic Variables. These 15 variables constitute the stressed subset from the full 132 FRED-MD and Goyal–Welch factors. This selection covers key transmission channels in financial stress testing: equity markets (S&P 500), inflation (CPI), foreign exchange (4 currency pairs vs. USD), monetary policy (Fed Funds rate, Treasury rates at 3-month, 5-year, and 10-year maturities), credit spreads (AAA and BAA corporate bonds), real economic activity (real personal income, civilian unemployment), and market volatility (VXO index).
@@ 7.2 Stress Test Analysis Data
We conduct a historical backtest as outlined in Algorithm 6 in Appendix F.5 for each one of the financial crises. For each backtest, in addition to reporting the performance metrics MAE and percentage accuracy, we also display the time series plots of the predicted conditional portfolio returns for each one of the models along with the true historical portfolio returns, where we re-emphasize that each portfolio is constructed with the $1/N$ rule. An important note is that for Static PCA, we perform each backtest for every possible PC component dimension and select the best performing one in terms of MAE. We begin with the richest of the four backtests, spanning June 2004-December 2016, where we have 150 prediction periods. As mentioned earlier, in this backtest we periodically refit the Kalman filters every 50 months. Figure 6 displays the time series of the conditional predictions from the backtest. We note that for a few months, Dynamic PCA yielded predictions that were significantly far off from the truth, we removed some of these outlier predictions from the plot to preserve the visibility of all the other time series displayed. We display a gray bar over each month where Dynamic PCA yielded such an outlier prediction.
Figure 6: Historical backtest: June 2004-December 2016 (150 out-of-sample monthly predictions). True portfolio returns (purple) versus conditional predictions from JDKF (red), SSA (blue), Static PCA (orange), and Dynamic PCA (green). Gray vertical bars indicate periods where Dynamic PCA produced outlier predictions. JDKF tracks realized returns more closely throughout, particularly during the 2007-2009 financial crisis. We consider an equal-weighted portfolio across 1,776 stocks. Scenarios are defined by stressing the 15 macroeconomic factors in Table LABEL:tab:scenario_variables.
@@
Table 4: Value-at-Risk exceptions test at 95% confidence level for JDKF predictions. Under the null hypothesis that JDKF is correct and its predicted portfolio returns are consistent with the realized returns, the number of exceptions (realized returns below the VaR threshold) should follow Binomial(150,0.05)(150,0.05) with expected value 7.5. We observed 7 exceptions. Two-sided binomial test yields $p=0.99$, providing no evidence against JDKF’s distributional accuracy. The VaR threshold of $6.91\%$ corresponds to the empirical 5th percentile of JDKF’s predicted conditional return distributions.
Note that for the 95% VaR exceptions test in Table 4 there is no strong evidence to reject the null hypothesis that JDKF is correct, i.e., that the VaR estimates from JDKF appear statistically consistent with the expected distribution. In both cases the number of observed exceptions is in line with the expected exceptions. For the 95% test we observe 7 exceptions with an expected 7.5, JDKF is slightly conservative but certainly well within an acceptable range. This indicates JDKF is not displaying systematic under or over-estimation of risk.
@@
Figure 7: Diffusion map diagnostics for selecting kernel bandwidth $\varepsilon$ and embedding dimension $\ell$ using the 132 macroeconomic factors. We select a bandwith parameter equal to the median $d(x(t_{i}),x(t_{j}))$.
@@
Table 5: Mean Absolute Error (MAE) of conditional portfolio return predictions across four historical periods. JDKF achieves the lowest MAE in all periods, with improvements of 35.86% (2004-2016), 55.43% (1990-1991), 12.75% (2001), and 39.17% (2007-2009) relative to the best-performing benchmark. Dynamic PCA performs poorly during crisis periods (1990-1991, 2007-2009) but remains competitive during the Dot-Com bubble (2001). Static PCA consistently outperforms SSA, demonstrating the value of capturing factor co-movements even without explicit temporal dynamics. All methods estimate factor loadings $\tilde{\mathbf{B}}$ via OLS on rolling windows. JDKF and Dynamic PCA additionally employ Kalman filtering in their respective embedding spaces (diffusion coordinates vs. principal components).
@@
Table 6: Percentage of prediction periods where JDKF achieves lower absolute error than each benchmark. Each cell shows $100\times\#\{t:E_{\text{JDKF},t}<E_{\text{benchmark},t}\}/T$ where $E_{t}=|V_{t}^{\text{true}}-V_{t}^{\text{pred}}|$ is the absolute prediction error at time $t$. JDKF outperforms benchmarks 62.5%-77.78% of the time across different crises. Perfect performance (100% vs. Static PCA in 2001) indicates JDKF was closer to realized returns in all 8 months of that period.
@@ 8 Conclusions
We combine anisotropic diffusion maps with Kalman filtering to infer the latent dynamic covariate embeddings, and predict the response variable directly from the diffusion map embedding space. We develop a conditional sampling procedure that accounts for dynamic correlations between covariates when conditioning on fixed values for a subset of them, and note that the procedure is general enough that it allows for conditioning on subsets of responses as well. Our approach is justified via theoretical results on the convergence of anisotropic diffusion maps as well as the robustness of approximating the learned diffusion coordinates by linear SDEs. The former result applies concentration inequalities for Markov processes to generalize the arguments of [38, 39] to the time series case; the latter adapts arguments typically seen in literature on Markov chain Monte Carlo methods.
@@
While the primary application of our framework in this paper has been financial stress testing, it is important to emphasize the rich set of applications where our framework can be employed. First, our dynamic factor modeling framework is well-suited for asset pricing problems, where one seeks to construct a sparse set of dynamic factors from a high-dimensional factor zoo that possess high explanatory power for the cross-section of stock returns. Another application of our framework is to the problem of nowcasting, where one seeks to utilize a higher-frequency dataset to improve predictions on lower-frequency data. Another application is nowcasting (see [20]), where one makes use of higher frequency/quality data to improve predictions on lower frequency/quality data. Our approach is well-suited for such a task as it produces lower-dimensional dynamic embeddings which propagate linearly, and also possess explanatory power comparable to the original high-dimensional set of factors.